OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 48π yd? A. d = 24 yd r = 12 yd B. d = 24 yd r = 48 yd C. d = 12 yd r = 24 yd D. d = 48 yd r = 24 yd id like to know how to get the answer as well!!!! :)

OpenStudy (tkhunny):

\(Circumference = \pi\cdot Diameter = \pi\cdot (2\cdot Radius)\) Go!

OpenStudy (anonymous):

thank you! :)